OpenStudy (anonymous):

Write an equation for each line in point slope form Slope = -6 ; through (-3, -2)

OpenStudy (gorv):

standdard equation of line y=mx+c here m is slope

OpenStudy (gorv):

replace slope by -6

OpenStudy (gorv):

plzzz rply

OpenStudy (anonymous):

Thank you . Is it y=-6x+-3/-2 ??

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

y=-6x+c -2=-3*(-6)+c -2-18=c c=-20

OpenStudy (gorv):

we have to replace x and y value not c

OpenStudy (anonymous):

Thank You very Much